find xy, x(-9,2) and y(5,-4)


1
Expert's answer
2021-03-07T09:50:21-0500

"xy=\sqrt{((5-(-9))^2+(-4-2)^2 )}=\sqrt{14^2+(-6)^2}=\sqrt{196+36}=\sqrt{232}=2\sqrt{58}"